# Generative Modeling through Spectral Analysis of Koopman Operator

### v1

@@ 2.2 Koopman Operator
$\mathcal{A}f(x)=\langle b(x),\nabla f(x)\rangle+\frac{1}{2}\text{Tr}\left(\sigma(x)\sigma(x)^{\top}\nabla^{2}f(x)\right),$ (2.7)
for all $x\in\mathcal{X}$ and $f\in C^{2}_{b}(\mathcal{X})\cap L^{2}(\mathcal{X})$.
In practice, the eigenpairs of the Koopman operator can be directly estimated from data. Several data-driven methods have been developed for Koopman operator approximation in deterministic or stochastic settings such as EDMD [Williams_2015] and etc. [arbabi2017ergodic, baikonode, cheng2025information, colbrook2024rigorous, gao2025conformalonlinelearningdeep, ishikawa2024koopman, kevrekidis2016kernel, kostic2022learning, mauroy2024analytic, oprea2025distributional, xu2025reinforced, 10.1063/5.0283640, xu2025reskoopnet, pmlr-v267-zhang25be].
3 Data-Driven Spectral Construction
3.1 Koopman Interpretation of Wasserstein Gradient Descent
While LAWGD [chewi2020svgd] provides a theoretically appealing sampling dynamics with scale-free exponential convergence, its formulation assumes access to the exact inverse of the Langevin generator. More precisely, for the overdamped Langevin dynamics, i.e., letting $b=-\nabla V$ and $\sigma=\sqrt{2}I_{d}$ for (2.5), the Langevin generator [pavliotis2016stochastic] is $\mathcal{L}=-\Delta+\langle\nabla V,\nabla\cdot\rangle$ [pavliotis2016stochastic]. In this idealized setting, the algorithm requires the full spectral decomposition of $\mathcal{L}$, and thus it is not directly implementable when $V$ or $\mathcal{L}$ are unknown.
A key observation is that the Langevin generator $\mathcal{L}$ coincides (up to a sign) with the Koopman generator of the Markov semigroup associated with $(X_{t})_{t\geq 0}$. In other words, if $(\mathcal{T}^{t})_{t\geq 0}$ denotes the Koopman semigroup, then for any smooth test function $f$, its infinitesimal generator $\mathcal{A}$ satisfies
$\mathcal{A}f(x)=-\langle\nabla V(x),\nabla f(x)\rangle+\Delta f(x)=-\mathcal{L}f(x),$
by (2.7). Hence $\mathcal{A}=$-$\mathcal{L}$ on the common domain of definition. This link makes Koopman spectral methods a natural tool for approximating the eigenpairs of $\mathcal{L}$ from data. Since LAWGD requires access to $\mathcal{L}^{-1}$, the Koopman spectral decomposition provides a natural finite-dimensional surrogate through the first $r$ eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$. We define the truncated inverse operator
$\mathcal{K}_{r}\coloneqq\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\,\langle\cdot,\phi_{i}\rangle_{\pi}\,\phi_{i},$ (3.1)
which recovers the exact inverse, i.e., $\mathcal{L}^{-1}$, as $r\to\infty$ as shown in [10.1063/5.0283640] and plays the role of the LAWGD kernel in the Koopman-based construction.
3.2 Truncated Koopman Approximation and Particle Dynamics
In practice, we approximate the leading eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$ of $\mathcal{L}$ through data-driven estimators of the Koopman generator $\mathcal{A}$, e.g., EDMD [Williams_2015] or kernel-EDMD [kevrekidis2016kernel]. Given training samples $\{z^{(j)}\}_{j=1}^{N}\sim\pi$, these methods construct a data-driven approximation $\widehat{\mathcal{A}}_{N}$ on a $N$-dimensional subspace whose leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ give the $r$-dimensional approximated inverse operator
$\widehat{\mathcal{K}}_{r}=\sum_{i=1}^{r}\frac{1}{\widehat{\lambda}_{i}}\langle\cdot,\widehat{\phi}_{i}\rangle_{\pi}\widehat{\phi}_{i}.$ (3.2)
This construction defines a Koopman-based approximation of the integral operator $\mathcal{K}_{\pi}$ constructed by the LAWGD kernel as in (2.2), preserving the functional form of the LAWGD dynamics while introducing spectral truncation from the finite rank $r$.
Next, the particle dynamics require evaluating the gradient of the truncated kernel
$K_{\widehat{\mathcal{K}}_{r}}(x,y)=\sum_{i=1}^{r}\frac{\widehat{\phi}_{i}(x)\widehat{\phi}_{i}(y)}{\widehat{\lambda}_{i}}.$
For each particle pair $(x^{(i)}_{t},x^{(j)}_{t})$, we compute
$\nabla_{1}K_{\widehat{\mathcal{K}}_{r}}(x^{(i)}_{t},x^{(j)}_{t})=\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x^{(i)}_{t})\cdot\widehat{\phi}_{k}(x^{(j)}_{t})}{\widehat{\lambda}_{k}},$ (3.3)
where $\nabla_{1}$ denotes the gradient with respect to the first argument and $\nabla\widehat{\phi}_{k}$ can be computed analytically for smooth basis functions (e.g., Gaussian kernels) or approximated via finite differences. Following (2.4), the Koopman-adjusted particle update becomes
$x_{t+1}^{(i)}=x_{t}^{(i)}-\frac{h}{M}\sum_{j=1}^{M}\nabla_{1}K_{\widehat{\mathcal{K}}_{r}}(x_{t}^{(i)},x_{t}^{(j)}),\quad 1\leq i\leq M.$ (3.4)
@@ 3.2 Truncated Koopman Approximation and Particle Dynamics
1
21exConstruct dictionary $\{\psi_{k}\}_{k=1}^{n}$ and estimate Koopman operator using trajectory pairs $\{(z^{(j)},z^{(j+1)})\}_{j=1}^{N-1}$ ;
3 Compute leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ of $\widehat{\mathcal{A}}_{N}$ ;
4
51exfor $t=0,1,\ldots,T-1$ do
6 for $i=1,\ldots,M$ do
7    $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$ ;
8    for $j=1,\ldots,M$ do
9       $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}+\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}$ by Eq. (3.3)
10    $x_{t+1}^{(i)}\leftarrow x_{t}^{(i)}-\frac{h}{M}\mathbf{v}_{i}$ by Eq. (3.4)
11
@@ 7    $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$ ;
Algorithm 1 KSWGD for Time Series Data
1
21exEstimate score: $\nabla\log\widehat{\pi}(x)$ using KDE ;
3 for $j=1,\ldots,N$ do
4   Sample $\xi^{(j)}\sim\mathcal{N}(0,I_{d})$ ;
5 $\hat{z}_{\Delta t}^{(j)}\leftarrow z^{(j)}-\Delta t\cdot\nabla\log\widehat{\pi}(z^{(j)})+\sqrt{2\Delta t}\,\xi^{(j)}$ ;
6
7
81exConstruct dictionary $\{\psi_{k}\}_{k=1}^{n}$ and estimate Koopman operator using pairs $\{(z^{(j)},\hat{z}_{\Delta t}^{(j)})\}_{j=1}^{N}$ ;
9 Compute leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ of $\widehat{\mathcal{A}}_{N}$ ;
10
111exfor $t=0,1,\ldots,T-1$ do
12 for $i=1,\ldots,M$ do
13    $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$ ;
14    for $j=1,\ldots,M$ do
15       $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}+\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}$
16    $x_{t+1}^{(i)}\leftarrow x_{t}^{(i)}-\frac{h}{M}\mathbf{v}_{i}$
17
Output: Generated particles $\{x_{T}^{(i)}\}_{i=1}^{M}$
Algorithm 2 KSWGD for Static Time Data
Algorithm 1 and  2 summarized the complete procedure for time series type of data and static time data (e.g., image dataset). The computational cost is dominated by the offline eigendecomposition ($O(n^{3})$ for basis size $n$) and the online kernel gradient evaluation ($O(M^{2}rd)$ per iteration for $M$ particles in dimension $d$ with rank $r$). Unlike score-based methods that require neural network training, KAWGD’s sampling phase is deterministic and training-free once the Koopman basis is determined. Notice that the gradient of eigenfunctions in (3.3) would be difficult to compute. In this work, we compute it by Diffusion map with Gaussian RBF kernel. See [li2025diffusion] for more details.
@@ 4.1 Data-Driven Error Bound Analysis
In this section, we analyze the convergence properties of the Koopman spectral Wasserstein gradient descent dynamics introduced in Section 3. Throughout the section, we denote
$\rho_{t}\coloneqq\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi},\qquad f_{t}\coloneqq\rho_{t}-1\in L_{0}^{2}(\pi),$
the fluctuation of the density with respect to $\pi$. Let $\{(\phi_{i},\lambda_{i})\}_{i\geq 1}$ be the eigenpairs of the Langevin generator $\mathcal{L}$, and recall the truncated inverse operator in (3.1), we have
$\mathcal{L}\mathcal{K}_{r}=\Pi_{r},$ (4.1)
@@
For regularity conditions, we assume that $\rho_{t}>0\textit{ a.e.}$ and $\partial_{t}(\rho_{t}\log\rho_{t})\in L^{1}_{\pi}$. In addition, we assume that $\nabla f_{t}\in L^{2}(\pi)$ and $\mathcal{K}_{r}f_{t}\in\mathcal{D}(\mathcal{L})$.
@@
Assume that Assumptions 4.1 and 4.2 holds. Let $(\mu_{t})_{t\geq 0}$ be the solution to the truncated LAWGD dynamics with exact eigenpairs and initial distribution $\mu_{0}$:
@@
where $\mathcal{K}_{r}=\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\langle\cdot,\phi_{i}\rangle_{\pi}\phi_{i}$ defined in (3.1) satisfies (4.1). Then
$\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)e^{-t}+\eta_{r}^{2}(1-e^{-t}).$ (4.2)

### v3

@@ 2.2 Koopman Operator
$\mathcal{A}f(x)=\langle b(x),\nabla f(x)\rangle+\frac{1}{2}\text{Tr}\left(\sigma(x)\sigma(x)^{\top}\nabla^{2}f(x)\right).$ (5)
This operator-theoretic perspective is powerful because the eigenpairs of $\mathcal{A}$ encode the global properties of the stochastic dynamics.
While $\mathcal{A}$ is infinite-dimensional, it can be approximated from finite trajectory data. Extended Dynamic Mode Decomposition (EDMD) (Williams et al., 2015) and its variants project the operator onto a finite-dimensional subspace. Recent advances extend this to reproducing kernel Hilbert spaces (Kevrekidis et al., 2016; Kostic et al., 2022; Ishikawa et al., 2024) and neural networks (Lusch et al., 2018; Bai and Ding, 2025; Zhang et al., 2025; Xu et al., 2025a), allowing for efficient spectral estimation even in high-dimensional stochastic settings (Arbabi and Mezic, 2017; Oprea et al., 2025; Xu et al., 2025b).
3 Koopman Spectral Wasserstein Gradient Descent
In this section, we present our main framework, Koopman Spectral Wasserstein Gradient Descent (KSWGD). The central goal is to construct a tractable approximation of the ideal preconditioner $\mathcal{K}_{\pi}=\mathcal{L}^{-1}$ solely from trajectory data, without explicit knowledge of the potential $V$. We achieve this by exploiting the spectral correspondence between the Langevin generator and the Koopman operator.
3.1 Operator-Theoretic Foundation
Let $\mathcal{L}=-\Delta+\langle\nabla V,\nabla\cdot\rangle$ be the Langevin generator on $L^{2}_{\pi}$. Assuming a discrete spectrum, the ideal preconditioner acts as
$\mathcal{K}_{\pi}f=\sum_{i=1}^{\infty}\lambda_{i}^{-1}\langle f,\phi_{i}\rangle_{\pi}\phi_{i},$ (6)
where $\{(\lambda_{i},\phi_{i})\}$ are the eigenpairs of $\mathcal{L}$ with $0=\lambda_{0}<\lambda_{1}\leq\dots$. Direct computation of Eq. (6) is, as mentioned in Section 2.1, generally impossible due to the unknown $V$.
However, consider the Koopman operator associated with the overdamped Langevin dynamics in Eq. (3). Its infinitesimal generator $\mathcal{A}$ satisfies
$\mathcal{A}f=-\langle\nabla V,\nabla f\rangle+\Delta f=-\mathcal{L}f,$
on their common domain. This identity $\mathcal{A}=$-$\mathcal{L}$ implies that the eigenfunctions of $\mathcal{A}$ are exactly those of $\mathcal{L}$, while the eigenvalues are simply negated. Consequently, we can approximate the spectral components of the inverse Langevin generator by estimating the spectrum of the Koopman generator from time-series data.
Defining the rank-$r$ truncation, we propose the spectral preconditioner:
$\mathcal{K}_{r}\coloneqq\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\,\langle\cdot,\phi_{i}\rangle_{\pi}\,\phi_{i},$ (7)
which approximates $\mathcal{L}^{-1}$ and recovers the exact inverse as $r\to\infty$ (Xu et al., 2025b). This truncated operator serves as our data-driven preconditioner in the Wasserstein gradient flow.
3.2 Data-Driven Spectral Approximation
To realize $\mathcal{K}_{r}$ in practice, we employ data-driven Koopman methods such as EDMD (Williams et al., 2015) or kernel-EDMD (Kevrekidis et al., 2016). Given training trajectories $\{(z^{(j)},z^{\prime(j)})\}_{j=1}^{N}$ where $z^{\prime(j)}$ is the state of $z^{(j)}$ after a short interval $\tau$, these methods approximate the generator $\mathcal{A}$ on a finite-dimensional subspace spanned by a dictionary (or kernel feature map).
Solving the generalized eigenvalue problem associated with the data matrices yields estimated eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$. Substituting these into (7) gives our empirical preconditioner:
$\widehat{\mathcal{K}}_{r}f(x)=\sum_{i=1}^{r}\frac{1}{\widehat{\lambda}_{i}}\langle f,\widehat{\phi}_{i}\rangle_{\pi}\widehat{\phi}_{i}(x).$ (8)
Crucially, this construction is training-free in the sense of deep learning; it requires only linear algebra operations and avoids iterative backpropagation.
3.3 Particle Dynamics and Algorithm
With the estimated spectral components, we formulate the particle update rule. The velocity field requires the gradient of the kernel $K_{\widehat{\mathcal{K}}_{r}}(x,y)=\sum_{i=1}^{r}\widehat{\lambda}_{i}^{-1}\widehat{\phi}_{i}(x)\widehat{\phi}_{i}(y)$ with respect to the first argument. Substituting this into the particle implementation (2), the update rule for the $i$-th particle $x_{t}^{(i)}$ is:
$x_{t+1}^{(i)}=x_{t}^{(i)}-\frac{h}{M}\sum_{j=1}^{M}\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}.$ (9)
Gradient Computation: A practical challenge is computing $\nabla\widehat{\phi}_{k}$ in (9), which is typically hard for general function approximators. In this work, we compute it analytically using a backward kernel matrix with a Gaussian RBF kernel (see (Li and Marzouk, 2025) for details). Algorithm 1 summarizes the complete procedure for time-series data. Importantly, our framework extends to static data (e.g., image datasets) by constructing synthetic dynamics; we provide the corresponding procedure in Algorithm 2 (Appendix A.6).
Complexity:
The computational cost is dominated by the offline eigendecomposition ($O(n^{3})$ for basis size $n$) and the online kernel gradient evaluation ($O(M^{2}rd)$ per iteration for $M$ particles, rank $r$, dimension $d$). Unlike score-based generative models (Song et al., 2020) that require extensive neural network training and stochastic sampling, KSWGD’s sampling phase is entirely deterministic and training-free once the Koopman basis is determined.
Algorithm 1 KSWGD for Time-Series Data
@@ 3.3 Particle Dynamics and Algorithm
 Construct dictionary $\{\psi_{k}\}_{k=1}^{n}$ and estimate the Koopman operator using trajectory pairs $\{(z^{(j)},z^{(j+1)})\}_{j=1}^{N-1}$.
 Compute leading $r$ eigenpairs $\{(\widehat{\lambda}_{k},\widehat{\phi}_{k})\}_{k=1}^{r}$ of $\widehat{\mathcal{A}}_{N}$.
for $t=0$ to T1T-1 do
  for $i=1$ to $M$ do
   Initialize $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$.
   for $j=1$ to $M$ do
    $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}+\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}.$ {see Eq. (9)}
   end for
   $x_{t+1}^{(i)}\leftarrow x_{t}^{(i)}-\frac{h}{M}\mathbf{v}_{i}.$ {see Eq. (9)}
  end for
end for
@@ 4 Convergence and Error Bound Analysis
This section develops the main theoretical guarantees for our proposed method KSWGD. We first analyze the continuous-time dynamics and derive error bounds that separate the ideal preconditioned flow from the effects of truncation and data-driven Koopman approximation in Section 4.1. We then extend these theory to the practical discrete-time algorithm via the Approximate Gradient Flow framework in Section 4.2, which shows a linear convergence (up to an explicit bias term) under suitable step-size conditions.
@@ 4.1 Data-Driven Error Bound Analysis
In this section, we analyze the convergence properties of the Koopman spectral Wasserstein gradient descent dynamics introduced in Section 3. Throughout the section, we denote $\rho_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}$ and
$f_{t}\coloneqq\rho_{t}-1\in L_{0}^{2}(\pi),$
the fluctuation of the density with respect to $\pi$. Let $\{(\phi_{i},\lambda_{i})\}_{i\geq 1}$ be the eigenpairs of the Langevin generator $\mathcal{L}$, and recall the truncated inverse operator in (7), we have
$\mathcal{L}\mathcal{K}_{r}=\Pi_{r},$ (10)
@@
For regularity conditions, we assume that $\rho_{t}>0\textit{ a.e.}$ and $\partial_{t}(\rho_{t}\log\rho_{t})\in L^{1}_{\pi}$. In addition, we assume that $\nabla f_{t}\in L^{2}_{\pi}$ and $\mathcal{K}_{r}f_{t}\in\mathcal{D}(\mathcal{L})$.
@@
Assume that Assumptions 4.1 and 4.2 hold. Let $(\mu_{t})_{t\geq 0}$ be the solution to the truncated LAWGD dynamics with exact eigenpairs and initial distribution $\mu_{0}$:
@@
where $\mathcal{K}_{r}=\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\langle\cdot,\phi_{i}\rangle_{\pi}\phi_{i}$ defined in (7) satisfies (10). Then
$\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)e^{-t}+\eta_{r}^{2}(1-e^{-t}).$ (11)